write an inequality to describe the situation and solve it to obtain the solution.
Your department sends its copying to the photocopying service down the street. The service bills your department $0.09 per page. Your department is considering purchasing a copier for $2750. The expense per page with the copier would be $0.04. If the expected life of the copier is 4 years how many copies must the department make in the 4-year period to justify the expenditure for the copier?

Answer by jainenderkapoor(61)


The department is paying $ 0.09 and it will be incurring a cost of $ 0.04 (in case the photocopier is purchased).
So the department will be saving = $0.09 - $0.04 = $0.05 per page
So the number of pages required to be copied so that the cost is equal to
cost of copier can be found by equation
$0.05 * x = 2750
which gives x = 2750/$0.05
x = 2750*100/5 = 55000 copies
so the department needs to make 55000 copies in four years to break-even (the cost of copier)

I hope that the solutin is clear to you.
If you have any doubt you are welcome to contact me.
I am also available for online tutoring in math at very reasonable rates.
My mail id is jainenderkapoor@gmail.com
Jainender Kapoor